OpenStudy (lemanual):

Can someone tell me What is the slope of the line which passes through (2, 5) and (3, 6)? (5 points) 2 0 3 1

pooja195 (pooja195):

Use the slope formula \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (lemanual):

Idk the formula My answer choice was 2 but I'm not really sure about it.

pooja195 (pooja195):

Well let's check the answer \[\LARGE \frac{6-5}{3-2}=slope\] Check your answer.

OpenStudy (igreen):

\(\sf (x_1, y_1), (x_2, y_2)\)

OpenStudy (lemanual):

So it's 1 then right?

OpenStudy (igreen):

Yes, that's correct.

OpenStudy (igreen):

You may give the medal to Pooja now :)

OpenStudy (lemanual):

Ty! Can you help me with one more please, If you have time? @iGreen

OpenStudy (igreen):

Sure thing

OpenStudy (lemanual):

Ok I'm gonna put up another question.